OpenStudy (itsmichelle29):

On the graph of the equation 6x − y = −1, what is the value of the y-intercept? -1 1 -6 6

OpenStudy (imtiaz7):

we use here slope intercept form that is y=mx+b

OpenStudy (itsmichelle29):

so its -1 ?

OpenStudy (imtiaz7):

6x-y=-1 y=mx+b -y=-6x-1 -y=-(6x+1) y=6x+1

OpenStudy (imtiaz7):

so now tell me the value of y intercept

OpenStudy (itsmichelle29):

1

OpenStudy (imtiaz7):

Good! :)